# Strong Linearizability without Compare&Swap: The Case of Bags

### v2

@@ 5.1 Linearizability
Next, suppose that $e$ is a read of Allocated on LABEL:WF-1-bounded:take:read_Allocated by a Take operation $tk$ into $a$, when Items[a] = $\bot$ or TS[a] = 1. Then $S^{\prime}$ is $S$ followed by $tk$ with return value EMPTY. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. $B$ is not empty if and only if there is a successful Insert operation in $S$ after which there is no successful Take operation. This correlates to $\sigma$ in which the producer wrote to Items on LABEL:WF-1-bounded:insert:write_item and there is no following successful t&s on TS. We will show that this is not the case in $S$. Assume, for sake of contradiction, that $\sigma$ contains a write of an element to Items by the producer on LABEL:WF-1-bounded:insert:write_item, but no successful t&s on TS after the last such write. Let $e$ be the step in which the producer executes this last write. By Lemma 4, TS[Allocated] = 0 at $e$. Since no successful t&s follows $e$ in $\sigma$, the producer may not pass LABEL:WF-1-bounded:insert:check_TS[m]. In particular, it may not reach LABEL:WF-1-bounded:insert:write_to_Allocated and thus Allocated remains unchanged. It also does not reach LABEL:WF-1-bounded:insert:clear_item, hence (and since Allocated remains unchanged), Items[Allocated] remains non-$\bot$ until the end of $\sigma$. The fact that Allocated remains unchanged also implies that TS[Allocated] remains 0 until the end of $\sigma$. But these contradict the fact that in the end of $\sigma$, Items[Allocated] = $\bot$ or TS[Allocated] = 1. Thus, $B$ is empty.
@@
Let $ins^{*}$ be the last successful Insert operation the producer performed before insins. Since $ins^{*}$ has completed, it is linearized before insins. Since $ins^{*}$ is linearized when it performs LABEL:WF-1-bounded:insert:write_item, the producer does not perform LABEL:WF-1-bounded:insert:clear_item through LABEL:WF-1-bounded:insert:write_item strictly between the linearization points of $ins^{*}$ and insins. Note that m is only changed on LABEL:WF-1-bounded:insert:choose_index and Allocated is only changed on LABEL:WF-1-bounded:insert:write_to_Allocated, so their values do not change during this interval. There are also no resets performed during this interval, since these only occur on LABEL:WF-1-bounded:insert:reset_TS. Moreover, by Lemma 1, throughout this interval, Items[Allocated] $\neq\bot$ and all other locations of Items contain $\bot$. Corollary 1 implies that the only successful t&s performed during this interval is on TS[Allocated]. However, if there is such a successful t&s, it is not reset before insins reads TS[m], so it would read 1, rather than 0. Hence, no successful Take operation is linearized between $ins^{*}$ and insins, and so $B$ contains the item that $ins^{*}$ inserted.
Finally, suppose $e$ is a write to Items by an Insert(x) operation insins on LABEL:WF-1-bounded:insert:write_item. Then $S^{\prime}$ is $S$ followed by insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. This is clearly true if there are no Insert operations in $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:WF-1-bounded:insert:write_item. By Lemma 4, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:WF-1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:WF-1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
@@ 22 Done: an ABA-detecting register
In this section, we present a lock-free, strongly-linearizable implementation of a bag that can contain at most one element. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). It combines ideas from our lock-free, strongly-linearizable implementation of an unbounded bag in Section 4 and our wait-free, linearizable implementation of a 1-bounded bag in Section 5. To keep the space bounded, we use an ABA-detecting register for Done, instead of a readable fetch&increment object. An ABA-detecting register has a lock-free, strongly-linearizable implementation from registers using bounded space [20].
@@ 6.1 Strong-Linearizability
Proof 6
All components of Items are initialized to $\bot$. The first successful Insert writes an element to Items[Allocated] on LABEL:1-bounded:insert:write_item. Any subsequent successful Insert writes $\bot$ to Items[Allocated] on LABEL:1-bounded:insert:clear_item. Note that the producer does not change the value of Allocated between LABEL:1-bounded:insert:write_item and LABEL:1-bounded:insert:clear_item, so, immediately after LABEL:1-bounded:insert:clear_item, all components of Items are $\bot$. Later, on LABEL:1-bounded:insert:write_item, the producer writes an element to Items[Allocated].
@@
Proof 7
The producer adds the location m = $k$ to used on LABEL:1-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:1-bounded:insert:clear_item. Thus, Items[k] = $\bot$ when $k$ is added to used. Suppose that an element is subsequently written to Items[k]. This occurs on LABEL:1-bounded:insert:write_item, with m = $k$. The value of m was chosen on LABEL:1-bounded:insert:choose_index to be a value that is not in hazardous. Thus, if $k$ was in used, it was removed from used on LABEL:1-bounded:insert:narrow_used_down, before an element was written to Items[m].
@@
Right after $D_{2}$, a t&s of TS[a] succeeds, so TS[a] = 0 at $D_{2}$. Assume, for sake of contradiction, that TS[a] = 1 at some point between $D_{0}$ and $D_{2}$. Then, after this point, but before $tk$ performed its successful t&s on TS[a], the producer must have reset TS[a] on LABEL:1-bounded:insert:reset_TS. By Lemma 6, Items[a] = $\bot$ when the reset was performed. Thus, between $D_{1}$ and $D_{2}$, the producer must have written $\bot$ to Items[a] on LABEL:1-bounded:insert:clear_item and then performed the reset. Between these two steps, the producer collects Hazards and reads $a$ from Hazards[i]. But, by LABEL:1-bounded:insert:reset_loop, the producer does not reset TS[a] on LABEL:1-bounded:insert:reset_TS. This is a contradiction. Thus TS[a] = 0 between $D_{0}$ and $D_{2}$.
After writing $x$ to Items[a] just before $D_{0}$, the producer next writes to Items[a] on LABEL:1-bounded:insert:clear_item. Between these two writes, the producer must read 1 from TS[a] on LABEL:1-bounded:insert:check_TS[m]. But TS[a] = 0 between $D_{0}$ and $D_{2}$. Thus, Items[a] = $x$ between $D_{0}$ and $D_{2}$.
@@
By Lemma 7, Items[a] = $\bot$ while $a$ was in used. Immediately after step $e$, Items[a] $\neq\bot$, so $a$ is not in used. Prior to being removed from used on LABEL:1-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:1-bounded:insert:reset_TS by some step $e^{\prime}$.
By Lemma 6, Items[a] = $\bot$ at $e^{\prime}$. Between $e^{\prime}$ and $e$, there was no write to Items[a] on LABEL:1-bounded:insert:write_item, so Items[a] = $\bot$. When a successful t&s on TS[a] is performed, Corollary 2 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s on TS[a] was performed between $e^{\prime}$ and $e$. This implies that TS[a] = 0 when $e$ was performed.
@@
First, suppose that $e$ is a successful t&s of TS[a] on LABEL:1-bounded:take:t&s by a Take operation, $tk$. Let $x\neq\bot$ be the value $tk$ read from Items[a] when it last performed LABEL:1-bounded:take:read_item during $\sigma$. Let $D_{1}$ be the configuration right before this read and let $D_{2}$ be the configuration at the end of $\sigma$. Let $D_{0}$ be the configuration right after the last write to Items[a] on LABEL:1-bounded:insert:write_item that was performed by the producer before $D_{1}$ and let insins be the Insert operation that performed this write. By Lemma 8, TS[a] = 0 and Items[a] = $x$ between $D_{0}$ and $D_{2}$. Therefore, by Lemma 6, no write to any cell of Items on LABEL:1-bounded:insert:write_item is performed between $D_{0}$ and $D_{2}$, which means insins performed the last write of an element to Items in $\sigma$.
@@
If there is a pending Insert(x’) operation, $ins^{\prime}$, at the end of $\sigma$ that wrote to Items[m] on LABEL:1-bounded:insert:write_item (but has not incremented Done on LABEL:1-bounded:insert:write_to_done), then $ins^{\prime}=ins$. In this case, $S^{\prime}$ is $S$ followed by insins with return value OK and then $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $B=\emptyset$. If there are no Insert operations in $S$, then the bag is empty at the end of $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:1-bounded:insert:write_item. By Lemma 9, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
@@
If there is no Insert operation that performed LABEL:1-bounded:insert:write_item in $\sigma$, but not LABEL:1-bounded:insert:write_to_done, then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $B=\{x\}$. Since insins performed LABEL:1-bounded:insert:write_item right before $D_{0}$, it also performed LABEL:1-bounded:insert:write_to_done and, thus, was completed in $\sigma$. Hence insins appears in $S$. Immediately after it appears, the bag contains only $x$. Between $D_{0}$ and $D_{2}$, TS[a] = 0, so no successful t&s is performed on TS[a]. Also, Items[a] $\neq\bot$, so, by Lemma 6, Items[a’] $\bot$ for all $a^{\prime}\neq a$. By Corollary 2, Items[a’] $\neq\bot$ when a successful t&s is performed on TS[a’]. Thus, no successful t&s is performed on any location of TS between $D_{0}$ and $D_{2}$. Hence, no Take operation that removes $x$ is linearized between $D_{0}$ and $D_{2}$. Since insins is linearized after $D_{0}$, there are no successful Take operations after insins in $S$. Therefore, at the end of $S$, the bag still contains $x$. Since $S$ is a linearization, it follows that $B=\{x\}$.
@@
So, suppose $B^{*}$ is not empty. Then it contains exactly one element, since $S$ is a linearization. Items inserted by coupled Insert operations are immediately taken from the bag by a Take operation linearized at the same step, so they are not in the bag in the end of the linearization of any prefix of the execution. Hence they are not in $B^{*}$. Therefore, the item, $x$, in $B^{*}$ was inserted into the bag by an uncoupled Insert operation, insins. This operation wrote $x$ to Items in some location, $a$, on LABEL:1-bounded:insert:write_item. When that occurred, TS[a] = 0 by Lemma 9. Since insins is uncoupled, no taker performs a successful t&s on TS[a] between this write and when insins writes to Done on LABEL:1-bounded:insert:write_to_done. Hence TS[a] = 0 at the linearization point of insins. No successful Insert follows insins in $S^{*}$, since $S^{*}$ is a linearization and the bag is 1-bounded. No successful Take operation follows insins in $S^{*}$, since it would remove $x$ from the bag. Thus, no successful t&s is performed between the linearization point of insins and $C^{*}$ and, so, TS[a] = 0 throughout this interval.
@@
Let $ins^{*}$ be the last successful Insert operation the producer performed before insins. Since $ins^{*}$ has completed, it is linearized before insins. Since $ins^{*}$ is linearized after it performs LABEL:1-bounded:insert:write_item, the producer does not perform LABEL:1-bounded:insert:clear_item through LABEL:1-bounded:insert:write_item between the linearization points of $ins^{*}$ and insins. Note that m is only changed on LABEL:1-bounded:insert:choose_index and Allocated is only changed on LABEL:1-bounded:insert:write_to_Allocated, so their values do not change during this interval. There are also no resets performed during this interval, since these only occur on LABEL:1-bounded:insert:reset_TS. Moreover, by Lemma 6, throughout this interval, Items[Allocated] $\neq\bot$ and all other locations of Items contain $\bot$. Corollary 2 implies that the only successful t&s performed during this interval is on TS[Allocated]. However, if there is such a successful t&s, it is not reset before insins reads TS[m], so it would read 1, rather than 0. Hence, no successful Take operation is linearized between $ins^{*}$ and insins, and so $B$ contains the item that $ins^{*}$ inserted.
Finally, suppose $e$ is a write to Done by an uncoupled Insert($x$) operation insins on LABEL:1-bounded:insert:write_to_done. Then $S^{\prime}$ is $S$ followed by insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. This is clearly true if there are no Insert operations in $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:1-bounded:insert:write_item. By Lemma 9, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.

### v3

@@ 5.1 Linearizability
Next, suppose that $e$ is a read of Allocated on LABEL:WF-1-bounded:take:read_Allocated by a Take operation $tk$ into $a$, when Items[a] = $\bot$ or TS[a] = 1. Then $S^{\prime}$ is $S$ followed by $tk$ with return value EMPTY. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. $B$ is not empty if and only if there is a successful Insert operation in $S$ after which there is no successful Take operation. This correlates to $\sigma$ in which the producer wrote to Items on LABEL:WF-1-bounded:insert:write_item and there is no following successful t&s on TS. We will show that this is not the case in $S$. Assume, for sake of contradiction, that $\sigma$ contains a write of an element to Items by the producer on LABEL:WF-1-bounded:insert:write_item, but no successful t&s on TS after the last such write. Let $e$ be the step in which the producer executes this last write. By Lemma 5.5, TS[Allocated] = 0 at $e$. Since no successful t&s follows $e$ in $\sigma$, the producer may not pass LABEL:WF-1-bounded:insert:check_TS[m]. In particular, it may not reach LABEL:WF-1-bounded:insert:write_to_Allocated and thus Allocated remains unchanged. It also does not reach LABEL:WF-1-bounded:insert:clear_item, hence (and since Allocated remains unchanged), Items[Allocated] remains non-$\bot$ until the end of $\sigma$. The fact that Allocated remains unchanged also implies that TS[Allocated] remains 0 until the end of $\sigma$. But these contradict the fact that in the end of $\sigma$, Items[Allocated] = $\bot$ or TS[Allocated] = 1. Thus, $B$ is empty.
@@
Let $ins^{*}$ be the last successful Insert operation the producer performed before insins. Since $ins^{*}$ has completed, it is linearized before insins. Since $ins^{*}$ is linearized when it performs LABEL:WF-1-bounded:insert:write_item, the producer does not perform LABEL:WF-1-bounded:insert:clear_item through LABEL:WF-1-bounded:insert:write_item strictly between the linearization points of $ins^{*}$ and insins. Note that m is only changed on LABEL:WF-1-bounded:insert:choose_index and Allocated is only changed on LABEL:WF-1-bounded:insert:write_to_Allocated, so their values do not change during this interval. There are also no resets performed during this interval, since these only occur on LABEL:WF-1-bounded:insert:reset_TS. Moreover, by Lemma 5.1, throughout this interval, Items[Allocated] $\neq\bot$ and all other locations of Items contain $\bot$. Corollary 5.4 implies that the only successful t&s performed during this interval is on TS[Allocated]. However, if there is such a successful t&s, it is not reset before insins reads TS[m], so it would read 1, rather than 0. Hence, no successful Take operation is linearized between $ins^{*}$ and insins, and so $B$ contains the item that $ins^{*}$ inserted.
Finally, suppose $e$ is a write to Items by an Insert(x) operation insins on LABEL:WF-1-bounded:insert:write_item. Then $S^{\prime}$ is $S$ followed by insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. This is clearly true if there are no Insert operations in $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:WF-1-bounded:insert:write_item. By Lemma 5.5, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:WF-1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:WF-1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
@@ 22 Done: an ABA-detecting register
In this section, we present a lock-free, strongly-linearizable implementation of a bag that can contain at most one element. It is shared by $n$ processes, $P_{1},\ldots,P_{n}$, called consumers, that can perform Take() and a single process, called the producer, that can perform Insert($x$). It combines ideas from our lock-free, strongly-linearizable implementation of an unbounded bag in Section 4 and our wait-free, linearizable implementation of a 1-bounded bag in Section 5. To keep the space bounded, we use an ABA-detecting register for Done, instead of a readable fetch&increment object. An ABA-detecting register has a lock-free, strongly-linearizable implementation from registers using bounded space [ovens2019strongly].
@@ 6.1 Strong-Linearizability
Proof.
All components of Items are initialized to $\bot$. The first successful Insert writes an element to Items[Allocated] on LABEL:1-bounded:insert:write_item. Any subsequent successful Insert writes $\bot$ to Items[Allocated] on LABEL:1-bounded:insert:clear_item. Note that the producer does not change the value of Allocated between LABEL:1-bounded:insert:write_item and LABEL:1-bounded:insert:clear_item, so, immediately after LABEL:1-bounded:insert:clear_item, all components of Items are $\bot$. Later, on LABEL:1-bounded:insert:write_item, the producer writes an element to Items[Allocated]. ∎
@@
Proof.
The producer adds the location m = $k$ to used on LABEL:1-bounded:insert:add_to_used, after it sets Items[m] to $\bot$ on LABEL:1-bounded:insert:clear_item. Thus, Items[k] = $\bot$ when $k$ is added to used. Suppose that an element is subsequently written to Items[k]. This occurs on LABEL:1-bounded:insert:write_item, with m = $k$. The value of m was chosen on LABEL:1-bounded:insert:choose_index to be a value that is not in hazardous. Thus, if $k$ was in used, it was removed from used on LABEL:1-bounded:insert:narrow_used_down, before an element was written to Items[m]. ∎
@@
Right after $D_{2}$, a t&s of TS[a] succeeds, so TS[a] = 0 at $D_{2}$. Assume, for sake of contradiction, that TS[a] = 1 at some point between $D_{0}$ and $D_{2}$. Then, after this point, but before $tk$ performed its successful t&s on TS[a], the producer must have reset TS[a] on LABEL:1-bounded:insert:reset_TS. By Lemma 6.1, Items[a] = $\bot$ when the reset was performed. Thus, between $D_{1}$ and $D_{2}$, the producer must have written $\bot$ to Items[a] on LABEL:1-bounded:insert:clear_item and then performed the reset. Between these two steps, the producer collects Hazards and reads $a$ from Hazards[i]. But, by LABEL:1-bounded:insert:reset_loop, the producer does not reset TS[a] on LABEL:1-bounded:insert:reset_TS. This is a contradiction. Thus TS[a] = 0 between $D_{0}$ and $D_{2}$.
After writing $x$ to Items[a] just before $D_{0}$, the producer next writes to Items[a] on LABEL:1-bounded:insert:clear_item. Between these two writes, the producer must read 1 from TS[a] on LABEL:1-bounded:insert:check_TS[m]. But TS[a] = 0 between $D_{0}$ and $D_{2}$. Thus, Items[a] = $x$ between $D_{0}$ and $D_{2}$. ∎
@@
By Lemma 6.2, Items[a] = $\bot$ while $a$ was in used. Immediately after step $e$, Items[a] $\neq\bot$, so $a$ is not in used. Prior to being removed from used on LABEL:1-bounded:insert:narrow_used_down, $a$ was not in hazardous, so TS[a] was reset to 0 on LABEL:1-bounded:insert:reset_TS by some step $e^{\prime}$.
By Lemma 6.1, Items[a] = $\bot$ at $e^{\prime}$. Between $e^{\prime}$ and $e$, there was no write to Items[a] on LABEL:1-bounded:insert:write_item, so Items[a] = $\bot$. When a successful t&s on TS[a] is performed, Corollary 6.4 implies that Items[a] $\neq$ $\bot$. Thus, no successful t&s on TS[a] was performed between $e^{\prime}$ and $e$. This implies that TS[a] = 0 when $e$ was performed. ∎
@@
First, suppose that $e$ is a successful t&s of TS[a] on LABEL:1-bounded:take:t&s by a Take operation, $tk$. Let $x\neq\bot$ be the value $tk$ read from Items[a] when it last performed LABEL:1-bounded:take:read_item during $\sigma$. Let $D_{1}$ be the configuration right before this read and let $D_{2}$ be the configuration at the end of $\sigma$. Let $D_{0}$ be the configuration right after the last write to Items[a] on LABEL:1-bounded:insert:write_item that was performed by the producer before $D_{1}$ and let insins be the Insert operation that performed this write. By Lemma 6.3, TS[a] = 0 and Items[a] = $x$ between $D_{0}$ and $D_{2}$. Therefore, by Lemma 6.1, no write to any cell of Items on LABEL:1-bounded:insert:write_item is performed between $D_{0}$ and $D_{2}$, which means insins performed the last write of an element to Items in $\sigma$.
@@
If there is a pending Insert(x’) operation, $ins^{\prime}$, at the end of $\sigma$ that wrote to Items[m] on LABEL:1-bounded:insert:write_item (but has not incremented Done on LABEL:1-bounded:insert:write_to_done), then $ins^{\prime}=ins$. In this case, $S^{\prime}$ is $S$ followed by insins with return value OK and then $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $B=\emptyset$. If there are no Insert operations in $S$, then the bag is empty at the end of $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:1-bounded:insert:write_item. By Lemma 6.5, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.
@@
If there is no Insert operation that performed LABEL:1-bounded:insert:write_item in $\sigma$, but not LABEL:1-bounded:insert:write_to_done, then $S^{\prime}$ is $S$ followed by $tk$ with return value $x$. To show that $S^{\prime}$ is a linearization, we must show that $B=\{x\}$. Since insins performed LABEL:1-bounded:insert:write_item right before $D_{0}$, it also performed LABEL:1-bounded:insert:write_to_done and, thus, was completed in $\sigma$. Hence insins appears in $S$. Immediately after it appears, the bag contains only $x$. Between $D_{0}$ and $D_{2}$, TS[a] = 0, so no successful t&s is performed on TS[a]. Also, Items[a] $\neq\bot$, so, by Lemma 6.1, Items[a’] $\bot$ for all $a^{\prime}\neq a$. By Corollary 6.4, Items[a’] $\neq\bot$ when a successful t&s is performed on TS[a’]. Thus, no successful t&s is performed on any location of TS between $D_{0}$ and $D_{2}$. Hence, no Take operation that removes $x$ is linearized between $D_{0}$ and $D_{2}$. Since insins is linearized after $D_{0}$, there are no successful Take operations after insins in $S$. Therefore, at the end of $S$, the bag still contains $x$. Since $S$ is a linearization, it follows that $B=\{x\}$.
@@
So, suppose $B^{*}$ is not empty. Then it contains exactly one element, since $S$ is a linearization. Items inserted by coupled Insert operations are immediately taken from the bag by a Take operation linearized at the same step, so they are not in the bag in the end of the linearization of any prefix of the execution. Hence they are not in $B^{*}$. Therefore, the item, $x$, in $B^{*}$ was inserted into the bag by an uncoupled Insert operation, insins. This operation wrote $x$ to Items in some location, $a$, on LABEL:1-bounded:insert:write_item. When that occurred, TS[a] = 0 by Lemma 6.5. Since insins is uncoupled, no taker performs a successful t&s on TS[a] between this write and when insins writes to Done on LABEL:1-bounded:insert:write_to_done. Hence TS[a] = 0 at the linearization point of insins. No successful Insert follows insins in $S^{*}$, since $S^{*}$ is a linearization and the bag is 1-bounded. No successful Take operation follows insins in $S^{*}$, since it would remove $x$ from the bag. Thus, no successful t&s is performed between the linearization point of insins and $C^{*}$ and, so, TS[a] = 0 throughout this interval.
@@
Let $ins^{*}$ be the last successful Insert operation the producer performed before insins. Since $ins^{*}$ has completed, it is linearized before insins. Since $ins^{*}$ is linearized after it performs LABEL:1-bounded:insert:write_item, the producer does not perform LABEL:1-bounded:insert:clear_item through LABEL:1-bounded:insert:write_item between the linearization points of $ins^{*}$ and insins. Note that m is only changed on LABEL:1-bounded:insert:choose_index and Allocated is only changed on LABEL:1-bounded:insert:write_to_Allocated, so their values do not change during this interval. There are also no resets performed during this interval, since these only occur on LABEL:1-bounded:insert:reset_TS. Moreover, by Lemma 6.1, throughout this interval, Items[Allocated] $\neq\bot$ and all other locations of Items contain $\bot$. Corollary 6.4 implies that the only successful t&s performed during this interval is on TS[Allocated]. However, if there is such a successful t&s, it is not reset before insins reads TS[m], so it would read 1, rather than 0. Hence, no successful Take operation is linearized between $ins^{*}$ and insins, and so $B$ contains the item that $ins^{*}$ inserted.
Finally, suppose $e$ is a write to Done by an uncoupled Insert($x$) operation insins on LABEL:1-bounded:insert:write_to_done. Then $S^{\prime}$ is $S$ followed by insins with return value OK. To show that $S^{\prime}$ is a linearization, we must show that $B$ is empty. This is clearly true if there are no Insert operations in $S$. Otherwise, let $ins^{*}$ be the last Insert operation in $S$. Let $e^{*}$ be the step of $\sigma$ in which $ins^{*}$ wrote to Items[$a^{*}$] on LABEL:1-bounded:insert:write_item. By Lemma 6.5, TS[$a^{*}$] = 0 when $e^{*}$ occurs. Since $S$ is a linearization, right after $ins^{*}$, the bag contains one item. Only successful Insert operations modify m (on LABEL:1-bounded:insert:choose_index), so m = $a^{*}$ when insins read 1 from TS[m] on LABEL:1-bounded:insert:check_TS[m]. Thus, a successful t&s on TS[$a^{*}$] must have been performed between $e^{*}$ and this step. The Take operation that performed it is linearized at this t&s. It appears after $ins^{*}$ in $S$, so the bag is empty at the end of $S$.